# What drives the growth of black holes: a decade of progress

### v1

@@ 4.1.3 Infrared selection
The dusty, molecular gas that surrounds accreting SMBHs will absorb optical/UV light, which is then re-emitted at IR wavelengths with the emission typically peaking at around 15–60 µm (Mullaney et al., 2011), reflecting the typical temperature of the dust (100 K). Detailed studies of nearby AGN systems over the last decade, exploiting the spatial information provided by mid-IR interferometry with the VLT (e.g. Hönig et al., 2013; Asmus et al., 2016; Asmus, 2019; Leftley et al., 2021), have shown that dust with a range of temperatures is typically present in a single AGN, with hotter dust likely associated with an outflowing wind extending in the polar direction, while cooler dust is likely associated with an inflowing, clumpy molecular gas disk that is also seen at sub-mm wavelengths (see §3.3 and Fig. 10 above).
@@ 4.1.4 Radio selection
It is difficult to distinguish between synchrotron emission from star formation and AGN activity. Brightness temperature measurements can pick out AGN cores independent of using a separate measure of radio excess, given high enough resolution (Condon et al., 1991). Using milli-arcsec resolution with wide-field VLBI techniques at GHz frequencies is now relatively developed (Garrett et al., 2001; Herrera Ruiz et al., 2017; Radcliffe et al., 2018), but still limited in field of view and number of sources, which makes it difficult to place these radio-selected AGN within the context of the general AGN population. Recent work using brightness temperature measurements from high-resolution LOFAR imaging, which increases field of view and sample sizes by over two orders of magnitude, has shown that even in non-radio excess sources, there can be a significant fraction (up to 50% more than expected) of radio emission contribution from AGN activity (Morabito et al., 2022b, 2025). This also becomes evident when forward-modelling the radio population in the LoTSS Deep Fields assuming a two-component model: star formation + AGN activity. Macfarlane et al. (2021) demonstrated that this two component model, where the AGN population is drawn from a power-law distribution down to low radio power, can completely reproduce the source counts and radio-loud fraction in the observations. What is also revealed is that there is an entire population of sources which would be classified as ‘radio-quiet’ but have more than half of their radio emission generated by AGN activity (Macfarlane et al., 2021, Fig. 21). The models are agnostic to whether the AGN radio emission is generated by a jet (as assumed by this single component model) as opposed to other radio AGN emitting processes, which allows for a new way to explore the radio properties of different AGN sub samples (Yue et al., 2024a).
@@ 4.2.1 Early theoretical and observational developments
4.2.2 Variability on Myr timescales explains the AGN–star formation connection
@@ 4.3 Jetted AGN – capping the growth of galaxies and their black holes
The substantial increase in the incidence of jetted AGN in galaxies with stellar mass persists even if an overall scaling of jet power with stellar mass is accounted for (e.g. Igo et al. 2024, although see also Hardcastle et al. 2019 for discussion of the impact of differing source lifetimes across the mass scale). The total kinetic power carried by jets, , may be estimated from the cavities they appear to inflate in the hot IGM of massive clusters or via assumptions about the synchrotron emitting plasma, as these are found to be in balance (Bîrzan et al., 2004; Rafferty et al., 2006; Timmerman et al., 2022). is found to scale (albeit with large scatter) with the jet radio luminosity212121The relation from a compilation of cavity-based measurements by Heckman and Best (2014), giving W, is widely adopted. The same quantity may be referred to as or as the “mechanical luminosity”, . and extrapolating this relationship allows estimates of the jet power more generally across the radio AGN population. We consider here the specific black hole kinetic power, sBHKP, where is the host galaxy stellar mass, as a mass-normalised tracer of the jet power in a manner analogous to the adoption of that provides a normalised measure of radiative power 222222As with (see Footnote 12), the specific black hole kinetic power can be given in “Eddington-ratio-equivalent” units if a single scaling relation between and is assumed, i.e. where is the kinetic power in erg s-1. Such units provide a useful scale to assess how much kinetic power is being released relative to the theoretical maximum radiative power from accretion onto a SMBH. (e.g. Igo et al., 2024). The top panel of Fig. 23 shows the distribution of for radio-selected AGN (i.e. sources where the radio emission is predominantly due to AGN) that are identified in deep LOFAR surveys (Kondapally et al., 2025), divided into those with higher radiative power (, where the radiative AGN luminosity, , has been estimated via SED fitting) and lower radiative power (), which should broadly correspond to radiatively efficient and inefficient phases.232323As discussed throughout this section, the imperfect correspondence between and is an important caveat when considering the true division of AGN based on their radiative power. Furthermore, recent studies indicate there is not a simple correspondence between higher and lower and the classification of sources as either HERGs or LERGs (e.g. Whittam et al., 2022). It is clear that jets spanning a broad range of are found across both AGN phases and weaker jets dominate in both cases. However, the highest jet power sources are associated with high radiative powers (i.e. high ). Thus, the power carried by jets appears to have a weak—but still significant—relationship to the current growth rate of the central SMBH.
@@ 4.4.1 Direct measurements of AGN incidence in groups and clusters
Over the past decade, direct measurements of the AGN populations in galaxy groups and clusters have extended to lower luminosities and higher redshifts thanks to improvements in observational sensitivity. For X-ray selected AGN, the AGN fraction in clusters (defined as the fraction of galaxies with AGN above some threshold) is found to decrease with increasing cluster mass, and decrease with radius toward the centers of clusters (e.g., Ehlert et al., 2014, 2015; Noordeh et al., 2020), with cluster outskirts reaching values approximately that of more typical “field” environments (e.g., Ehlert et al., 2014; Koulouridis et al., 2024). The X-ray AGN population in clusters also evolves rapidly with redshift, with the AGN fraction increasing by more than an order of magnitude to 1.5 (e.g., Martini et al., 2013), and optical- and IR-selected AGN in clusters show similar behavior (e.g., Alberts et al., 2016; Lopes et al., 2017). There has been some evidence presented for a potential excess of AGN outside the cluster virial radius, indicating fuelling of SMBH growth during galaxy infall (e.g., Fassbender et al., 2012; Koulouridis and Bartalucci, 2019), but these effects are not seen in all studies and may simply be owing to statistical fluctuations and projection effects (e.g., Muñoz Rodríguez et al., 2024).
@@ 4.5.2 Overview of results for merger fuelling
Using the technique (a) above, a broad range of studies collectively find no clear difference in merger fractions between AGN and “inactive” galaxies that are matched in key properties, leading to the conclusion that the bulk of AGN activity is not associated with mergers (e.g., Cisternas et al., 2011b; Villforth, 2023). There are some hints that mergers become more prevalent at the highest luminosities or for dust-reddened systems (e.g., Treister et al., 2012; Glikman et al., 2015) although these results have been contested (e.g., Villforth, 2023).
@@ 4.6 Black holes and AGN in dwarf galaxies
Currently, there is an ongoing systematic search for the presence of BHs in dwarf galaxies in the local Universe (Greene et al., 2008; Reines and Volonteri, 2015; Reines and Comastri, 2016; Mezcua et al., 2016; Mezcua, 2017; Chilingarian et al., 2018; Greene et al., 2020; Reines et al., 2020). The least massive BHs were found in RGG118, with a mass of (Baldassare et al., 2015), and in NGC4395 with a mass estimate of (Woo et al., 2019). The identification and confirmation of these low-mass BHs rely on multi-wavelength observations including optical spectroscopy, observations in X-ray, radio, and IR (e.g., Greene and Ho, 2004; Greene et al., 2008; Dong et al., 2012; Chilingarian et al., 2018; Liu et al., 2018b; Mezcua et al., 2018, 2019; Kaviraj et al., 2019), and remain challenging due to potential contamination from various sources, such as young starburst regions in low-metallicity environment (reassembling AGN), winds from both supernovae (SN) and massive stars, and emission from X-ray binaries (e.g., Gallo et al., 2008, 2010; Reines et al., 2011; Mezcua et al., 2016; Pardo et al., 2016; She et al., 2017; Mezcua et al., 2018; Fornasini et al., 2018; Schirra et al., 2020). Recent observations revealed that, in the low-mass regime, the presence of BHs is not confined to the centres of galaxies; e.g., Reines et al. (2020) discovered that half a dozen dwarf galaxies were hosting offset radio sources that could be AGN, and faint off-centre AGN signatures were detected in dwarf galaxies through MaNGA IFU observations (Mezcua and Domínguez Sánchez, 2020).
Most constraints indicate an AGN occupation fraction in dwarf galaxies, across all luminosity ranges, of , with significant variations spanning more than an order of magnitude (Shi et al., 2008; Aird et al., 2012; Schramm and Silverman, 2013; Reines et al., 2013; Miller et al., 2015; Lemons et al., 2015; Pardo et al., 2016; Aird et al., 2018; Mezcua et al., 2018; Kaviraj et al., 2019; Mezcua and Domínguez Sánchez, 2020; Birchall et al., 2020, 2022). Several studies have derived constraints on the BH occupation fraction, requiring corrections for inactive BHs missed in observational samples. These studies suggest that over half of galaxies with stellar masses of may harbor a central BH (Miller et al., 2015; Trump et al., 2015; Nguyen et al., 2019). In large-scale cosmological simulations, the BH occupation fraction is of almost unity in low-redshift low-mass galaxies. However, results for the AGN occupation fraction vary significantly, spanning several orders of magnitude due to differences in the subgrid modeling of BH seeding, growth, feedback and galaxy processes among simulations (Haidar et al., 2022).
@@ 5.3 Evidence for a quasar evolutionary sequence
Numerical simulations predict that a large fraction of the obscuration towards luminous AGN and quasars is due to intense on-going dust-obscured star formation, which can be enhanced in gas-rich galaxy mergers (e.g., Hopkins et al., 2008; Blecha et al., 2018). Observationally, it has been shown that the host galaxy can contribute to a substantial fraction of the obscuration towards the AGN (e.g., Buchner and Bauer, 2017; Circosta et al., 2019; Gilli et al., 2022; Alonso-Tetilla et al., 2024). Focusing on mid-IR selected quasars, Chen et al. (2015) and Andonie et al. (2022, 2024) showed that the fraction of obscured quasars increases with SFR such that the majority of quasars are obscured at the highest SFRs ( yr-1). Using high-resolution ALMA observations, Andonie et al. (2024) found that the star formation at the highest SFRs is so compact ( 0.5–3 kpc radius) that dust within the host galaxy can obscure the quasar even up to Compton-thick column densities ( cm-2). There is also evidence that the dust-covering factor in the nuclear regions is larger in the most heavily obscured systems, potentially due to chaotic gas inflows and/or radiation-pressure driven dusty outflows (e.g., Ramos Almeida et al., 2011; Yamada et al., 2021, 2024), and some obscured quasars may be completely buried so that even their AGN-powered narrow-emission lines are weak or undetected (e.g., Greenwell et al., 2022, 2024a). At the extreme end of the obscured quasar population are Hot Dust Obscured Galaxies (hot DOGs; see Table 2), which are amongst the most luminous AGN known and are heavily obscured with large dust covering fractors (e.g., Wu et al., 2012; Stern et al., 2014; Assef et al., 2015; Fan et al., 2016b). They appear to be growing close to their maximum rates but are not exclusively merger driven (Fan et al., 2016a; Wu et al., 2018; Díaz-Santos et al., 2021).
@@
The majority of red quasar studies over the last decade have shown that they have the properties expected for a transition population (prevalance for galaxy mergers, accreting at high Eddington ratios, and driving powerful outflows; e.g., Urrutia et al., 2012; Glikman et al., 2012, 2015; Banerji et al., 2015; DiPompeo et al., 2018; Calistro Rivera et al., 2021; Stacey et al., 2022; Kim et al., 2024a). ERQs appear to represent the extreme end of the red quasar distribution with the largest luminosities, highest-velocity outflows, and absorbing column densities (e.g., Zakamska et al., 2016; Perrotta et al., 2019; Ma et al., 2024; Vayner et al., 2024). However, depending on the selection of the “control” sample, normal blue quasars can also be identified with similar properties to those of red quasars (e.g., Zakamska et al., 2019; Villar Martín et al., 2020).
@@ 5.4 The high-redshift frontier of quasars
Fig. 29 charts the history of record-breaking quasar redshifts from the first identified quasar (Schmidt, 1963) to the present day (cf. Stark et al., 2025, for record-breaking galaxy redshifts). The 5 barrier was broken in 1999 (Fan et al., 1999) and the first quasar was identified in 2011 (Mortlock et al., 2011), requiring near-IR photometry; the highest-redshift quasar lies at (Wang et al., 2021b). JWST has pushed the identification of potential AGN to 10 with sensitive near–IR spectroscopy (Goulding et al., 2023; Bogdán et al., 2024; Maiolino et al., 2024b; Napolitano et al., 2024). See Fig. 26 for spectra of the three 10 systems, all of which have relatively narrow emission lines (FWHM1000 km s-1). The highest-redshift JWST AGN with broad Balmer emission lines (FWHM1000 km s-1) lie at 8.500 (Kokorev et al., 2023), 8.632 (Tripodi et al., 2024), (Larson et al., 2023), and (Taylor et al., 2025).

### v2

@@ 4.1.3 Infrared selection
The dusty, molecular gas that surrounds accreting SMBHs will absorb optical/UV light, which is then re-emitted at IR wavelengths with the emission typically peaking at around 15–60 µm (Mullaney et al., 2011), reflecting the typical temperature of the dust (100 K). Detailed studies of nearby AGN systems over the last decade, exploiting the spatial information provided by high-resolution mid-IR observations using large ground-based 8–10m telescopes (e.g. Hönig et al., 2013; Asmus et al., 2016; Asmus, 2019; Leftley et al., 2021; Garc´ıa-Bernete et al., 2022), have shown that dust with a range of temperatures is typically present in a single AGN, with hotter dust likely associated with an outflowing wind extending in the polar direction, while cooler dust is likely associated with an inflowing, clumpy molecular gas disk that is also seen at sub-mm wavelengths (see §3.3 and Fig. 10 above).
@@ 4.1.4 Radio selection
It is difficult to distinguish between synchrotron emission from star formation and AGN activity. Brightness temperature measurements can pick out AGN cores independent of using a separate measure of radio excess, given high enough resolution (Condon et al., 1991). Using milli-arcsec resolution with wide-field VLBI techniques at GHz frequencies is now relatively developed (Garrett et al., 2001; Herrera Ruiz et al., 2017; Radcliffe et al., 2018), but still limited in field of view and number of sources, which makes it difficult to place these radio-selected AGN within the context of the general AGN population. Recent work using brightness temperature measurements from high-resolution LOFAR imaging, which increases field of view and sample sizes by over two orders of magnitude, has shown that even in non-radio excess sources, there can be a significant fraction (up to 50% more than expected) of radio emission contribution from AGN activity (Morabito et al., 2022b, 2025). This also becomes evident when forward-modelling the radio population in the LoTSS Deep Fields assuming a two-component model: star formation + AGN activity. Macfarlane et al. (2021) demonstrated that this two component model, where the AGN population is drawn from a power-law distribution down to low radio power, can completely reproduce the source counts and radio-loud fraction in the observations. What is also revealed is that there is an entire population of sources which would be classified as ‘radio-quiet’ but have more than half of their radio emission generated by AGN activity (White et al., 2017; Macfarlane et al., 2021, see their Fig. 12). The models are agnostic to whether the AGN radio emission is generated by a jet (as assumed by this single component model) as opposed to other radio AGN emitting processes, which allows for a new way to explore the radio properties of different AGN sub samples (Yue et al., 2024a).
@@ 4.2.1 Early theoretical and observational developments
4.2.2 Variability on mega-year timescales explains the AGN–star formation connection
@@ 4.3 Jetted AGN – capping the growth of galaxies and their black holes
The substantial increase in the incidence of jetted AGN in galaxies with stellar mass persists even if an overall scaling of jet power with stellar mass is accounted for (e.g. Igo et al. 2024, although see also Hardcastle et al. 2019 for discussion of the impact of differing source lifetimes across the mass scale). The total kinetic power carried by jets, , may be estimated from the cavities they appear to inflate in the hot IGM of massive clusters or via assumptions about the synchrotron emitting plasma, as these are found to be in balance (Bˆırzan et al., 2004; Rafferty et al., 2006; Timmerman et al., 2022). is found to scale (albeit with large scatter) with the jet radio luminosity212121The relation from a compilation of cavity-based measurements by Heckman and Best (2014), giving W, is widely adopted. The same quantity may be referred to as or as the “mechanical luminosity”, . and extrapolating this relationship allows estimates of the jet power more generally across the radio AGN population. We consider here the specific black hole kinetic power, sBHKP, where is the host galaxy stellar mass, as a mass-normalised tracer of the jet power in a manner analogous to the adoption of that provides a normalised measure of radiative power 222222As with (see Footnote 12), the specific black hole kinetic power can be given in “Eddington-ratio-equivalent” units if a single scaling relation between and is assumed, i.e. where is the kinetic power in erg s-1. Such units provide a useful scale to assess how much kinetic power is being released relative to the theoretical maximum radiative power from accretion onto a SMBH. (e.g. Igo et al., 2024). The top panel of Fig. 23 shows the distribution of for radio-selected AGN (i.e. sources where the radio emission is predominantly due to AGN) that are identified in deep LOFAR surveys (Kondapally et al., 2025), divided into those with higher radiative power (, where the radiative AGN luminosity, , has been estimated via SED fitting) and lower radiative power (), which should broadly correspond to radiatively efficient and inefficient phases.232323As discussed throughout this section, the imperfect correspondence between and is an important caveat when considering the true division of AGN based on their radiative power. Furthermore, recent studies indicate there is not a simple correspondence between higher and lower and the classification of sources as either HERGs or LERGs (e.g. Whittam et al., 2022). It is clear that jets spanning a broad range of are found across both AGN phases and weaker jets dominate in both cases. However, the highest jet power sources are associated with high radiative powers (i.e. high ). Thus, the power carried by jets appears to have a weak—but still significant—relationship to the current growth rate of the central SMBH.
@@ 4.4.1 Direct measurements of AGN incidence in groups and clusters
Over the past decade, direct measurements of the AGN populations in galaxy groups and clusters have extended to lower luminosities and higher redshifts thanks to improvements in observational sensitivity. For X-ray selected AGN, the AGN fraction in clusters (defined as the fraction of galaxies with AGN above some threshold) is found to decrease with increasing cluster mass, and decrease with radius toward the centers of clusters (e.g., Ehlert et al., 2014, 2015; Noordeh et al., 2020), with cluster outskirts reaching values approximately that of more typical “field” environments (e.g., Ehlert et al., 2014; Koulouridis et al., 2024). The X-ray AGN population in clusters also evolves rapidly with redshift, with the AGN fraction increasing by more than an order of magnitude to 1.5 (e.g., Martini et al., 2013), and optical- and IR-selected AGN in clusters show similar behavior (e.g., Alberts et al., 2016; Lopes et al., 2017). There has been some evidence presented for a potential excess of AGN outside the cluster virial radius, indicating fuelling of SMBH growth during galaxy infall (e.g., Fassbender et al., 2012; Koulouridis and Bartalucci, 2019), but these effects are not seen in all studies and may simply be owing to statistical fluctuations and projection effects (e.g., Muñoz Rodr´ıguez et al., 2024).
@@ 4.5.2 Overview of results for merger fuelling
Using the technique (a) above, a broad range of studies collectively find no clear difference in merger fractions between AGN and “inactive” galaxies that are matched in key properties, leading to the conclusion that the bulk of AGN activity is not associated with mergers (e.g., Cisternas et al., 2011b; Villforth, 2023). There is evidence that mergers become more prevalent at the highest luminosities or for dust-reddened systems (e.g., Treister et al., 2012; Glikman et al., 2015) although these results have been contested (e.g., Villforth, 2023).
@@ 4.6 Black holes and AGN in dwarf galaxies
Currently, there is an ongoing systematic search for the presence of BHs in dwarf galaxies in the local Universe (Greene et al., 2008; Reines and Volonteri, 2015; Reines and Comastri, 2016; Mezcua et al., 2016; Mezcua, 2017; Chilingarian et al., 2018; Greene et al., 2020; Reines et al., 2020). The least massive BHs were found in RGG118, with a mass of (Baldassare et al., 2015), and in NGC4395 with a mass estimate of (Woo et al., 2019). The identification and confirmation of these low-mass BHs rely on multi-wavelength observations including optical spectroscopy, observations in X-ray, radio, and IR (e.g., Greene and Ho, 2004; Greene et al., 2008; Dong et al., 2012; Chilingarian et al., 2018; Liu et al., 2018b; Mezcua et al., 2018, 2019; Kaviraj et al., 2019), and remain challenging due to potential contamination from various sources, such as young starburst regions in low-metallicity environment (reassembling AGN), winds from both supernovae (SN) and massive stars, and emission from X-ray binaries (e.g., Gallo et al., 2008, 2010; Reines et al., 2011; Mezcua et al., 2016; Pardo et al., 2016; She et al., 2017; Mezcua et al., 2018; Fornasini et al., 2018; Schirra et al., 2020). Recent observations revealed that, in the low-mass regime, the presence of BHs is not confined to the centres of galaxies; e.g., Reines et al. (2020) discovered that half a dozen dwarf galaxies were hosting offset radio sources that could be AGN, and faint off-centre AGN signatures were detected in dwarf galaxies through MaNGA IFU observations (Mezcua and Dom´ınguez Sánchez, 2020).
Most constraints indicate an AGN occupation fraction in dwarf galaxies, across all luminosity ranges, of , with significant variations spanning more than an order of magnitude (Shi et al., 2008; Aird et al., 2012; Schramm and Silverman, 2013; Reines et al., 2013; Miller et al., 2015; Lemons et al., 2015; Pardo et al., 2016; Aird et al., 2018; Mezcua et al., 2018; Kaviraj et al., 2019; Mezcua and Dom´ınguez Sánchez, 2020; Birchall et al., 2020, 2022). Several studies have derived constraints on the BH occupation fraction, requiring corrections for inactive BHs missed in observational samples. These studies suggest that over half of galaxies with stellar masses of may harbor a central BH (Miller et al., 2015; Trump et al., 2015; Nguyen et al., 2019). In large-scale cosmological simulations, the BH occupation fraction is of almost unity in low-redshift low-mass galaxies. However, results for the AGN occupation fraction vary significantly, spanning several orders of magnitude due to differences in the subgrid modeling of BH seeding, growth, feedback and galaxy processes among simulations (Haidar et al., 2022).
@@ 5.3 Evidence for a quasar evolutionary sequence
Numerical simulations predict that a large fraction of the obscuration towards luminous AGN and quasars is due to intense on-going dust-obscured star formation, which can be enhanced in gas-rich galaxy mergers (e.g., Hopkins et al., 2008; Blecha et al., 2018). Observationally, it has been shown that the host galaxy can contribute to a substantial fraction of the obscuration towards the AGN (e.g., Buchner and Bauer, 2017; Circosta et al., 2019; Gilli et al., 2022; Alonso-Tetilla et al., 2024). Focusing on mid-IR selected quasars, Chen et al. (2015) and Andonie et al. (2022, 2024) showed that the fraction of obscured quasars increases with SFR such that the majority of quasars are obscured at the highest SFRs ( yr-1). Using high-resolution ALMA observations, Andonie et al. (2024) found that the star formation at the highest SFRs is so compact ( 0.5–3 kpc radius) that dust within the host galaxy can obscure the quasar even up to Compton-thick column densities ( cm-2). There is also evidence that the dust-covering factor in the nuclear regions is larger in the most heavily obscured systems, potentially due to chaotic gas inflows and/or radiation-pressure driven dusty outflows (e.g., Ramos Almeida et al., 2011; Yamada et al., 2021, 2024), and some obscured quasars may be completely buried so that even their AGN-powered narrow-emission lines are weak or undetected (e.g., Greenwell et al., 2022, 2024a). At the extreme end of the obscured quasar population are Hot Dust Obscured Galaxies (hot DOGs; see Table 2), which are amongst the most luminous AGN known and are heavily obscured with large dust covering fractors (e.g., Wu et al., 2012; Stern et al., 2014; Assef et al., 2015; Fan et al., 2016b). They appear to be growing close to their maximum rates but are not exclusively merger driven (Fan et al., 2016a; Wu et al., 2018; D´ıaz-Santos et al., 2021).
@@
The majority of red quasar studies over the last decade have shown that they have the properties expected for a transition population (prevalance for galaxy mergers, accreting at high Eddington ratios, and driving powerful outflows; e.g., Urrutia et al., 2012; Glikman et al., 2012, 2015; Banerji et al., 2015; DiPompeo et al., 2018; Calistro Rivera et al., 2021; Stacey et al., 2022; Kim et al., 2024a). ERQs appear to represent the extreme end of the red quasar distribution with the largest luminosities, highest-velocity outflows, and absorbing column densities (e.g., Zakamska et al., 2016; Perrotta et al., 2019; Ma et al., 2024; Vayner et al., 2024). However, depending on the selection of the “control” sample, normal blue quasars can also be identified with similar properties to those of red quasars (e.g., Zakamska et al., 2019; Villar Mart´ın et al., 2020).
@@ 5.4 The high-redshift frontier of quasars
Fig. 29 charts the history of record-breaking quasar redshifts from the first identified quasar (Schmidt, 1963) to the present day (cf. Stark et al., 2025, for record-breaking galaxy redshifts). The 5 barrier was broken in 1999 (Fan et al., 1999) and the first quasar was identified in 2011 (Mortlock et al., 2011), requiring near-IR photometry; the highest-redshift quasar lies at (Wang et al., 2021b). JWST has pushed the identification of potential AGN to 10 with sensitive near–IR spectroscopy (Goulding et al., 2023; Bogdán et al., 2024; Maiolino et al., 2024b; Napolitano et al., 2024). See Fig. 26 for spectra of the three 10 systems, all of which have relatively narrow emission lines (FWHM1000 km s-1). The highest-redshift JWST AGN with clear broad Balmer emission lines (FWHM1000 km s-1) lie at 8.500 (Kokorev et al., 2023), 8.632 (Tripodi et al., 2024), (Larson et al., 2023), and (Taylor et al., 2025).